# An Inversion-based Measure of Memorization for Diffusion Models

### v2

@@ 4.4 Text-to-image DM
For text-to-image DM, we model the prompt part as a distribution as well and optimize it simultaneously with the noise distribution. Suppose each input prompt is tokenized into maximum $N$ tokens $\bm{\omega}=[\bm{\omega}_{1},\bm{\omega}_{2},...,\bm{\omega}_{N}]$. Motivated by Guo et al. [20], we learn a $|\mathcal{V}|$-dimensional categorical distribution over the vocabulary $\mathcal{V}$ for each token $\bm{\omega}_{i}$. Let token $\bm{\omega}_{i}$ follows a categorical distribution with probabilities $\bm{\pi}_{i,j}$, where $\sum_{j=1}^{|\mathcal{V}|}\bm{\pi}_{i,j}=1$. $\bm{\pi}_{i,j}$ can be parameterized and optimized.
Because sampling from categorical distribution is non-differentiable, we adopt the Gumbel-Softmax reparameterization [23] to obtain smoothed sample $\tilde{\bm{\omega}}_{i}$. Given the smoothed token $\tilde{\bm{\omega}}_{i}$, it can be converted to token embedding as a linear combination of vocabulary embeddings $e(\tilde{\bm{\omega}}_{i})=\sum_{j=1}^{|\mathcal{V}|}\tilde{\bm{\omega}}_{i,j}e(j)$, where $e(j)$ is the pretrained token embedding of the $j$-th token. A proper choice of the prompt distribution is critical for inversion on text-to-image DM. We provide a comprehensive analysis in the supplementary material.
@@ 5 Experiment
In this section, we conduct a series of experiments from different perspectives to demonstrate the validity of InvMM. First, we conduct controlled studies to verify the effectiveness of the adaptive algorithm (Sec. 5.2). Based on this, we show that InvMM can accurately detect highly memorized images (Sec. 5.3) and elucidate the influence of various factors on memorization (Sec. 5.4). Finally, we discuss how memorization differs from membership (Sec. 5.5).
@@ 5.1 Experiment setup
Dataset. We use four datasets for evaluation, i.e., CIFAR-10 [30], CelebAHQ [25], FFHQ [26] and LAION-Aesthetics V2 [45]. For subsequent experiments, we construct four randomly sampled subsets of CelebAHQ and FFHQ, CelebAHQ-250, CelebAHQ-2.5k, FFHQ-600 and FFHQ-6k. We ensure that the larger subset includes the smaller subset. The LAION dataset is extremely large. To keep computational costs manageable, we collect a subset of images. We first include 76 heavily memorized images discovered by Webster [57]. We call this set of images the confirmed set. Furthermore, we collected another 50 images with high duplication rates according to the statistics from Webster et al [59]. This set of suspects, named as the suspicious set, are assured to be disjoint with the confirmed set. Images in the confirmed set can be triggered by their training captions while those in the suspicous set cannot. Finally, another 50 randomly sampled images are used for comparison, referred to as the normal set.
Diffusion Models. Following the implementations of DDPM [22] and Latent Diffusion Model (LDM) [41], we train DDPM on CIFAR-10 and LDM on the subsets of CelebAHQ and FFHQ. Pre-trained LDMs on the whole CelebAHQ and FFHQ are also used. For text-to-image generation, we utilize Stable Diffusion (SD) v1.4 and v2.1 trained on LAION.
Inversion. The experiment setting and additional results, including prompt inversion for SD, are deferred to the supplementary material.
@@ 5.2 Adaptive algorithm
A comparison of the resulting memorization score under different settings can be observed in Fig. 3. Fixing $\lambda$ to 1 cannot successfully invert most images: 92.4% and 99.7% images of CelebAHQ and FFHQ are assigned infinite memorization scores. In contrast, both $\lambda=0$ and adaptive reweighting successfully invert all images. However, Fixing $\lambda$ to 0 leads to larger resulting memorization scores, which indicates underestimation of the extent of memorization.
Figure 4 illustrates the generation results with the learned distributions. With the adaptive algorithm, we are able to successfully perform inversion on various datasets and DMs. The random generated samples closely resembles the training images. The results support the premise of InvMM being a meaningful measure.
@@ 5.3 Detection
$\bm{\epsilon}_{0}$-loss 0.981 0.414 1.000 1.000
$\bm{x}_{0}$-loss 0.180 0.000 0.992 0.961
@@
Wen [61] - - 1.000 1.000
InvMM* - - 1.000 1.000
InvMM 0.997 0.899 0.994 0.842
Table 1: Detection performance on DDPM and SD v1.4. *: using training captions, fixed during inversion.
@@
For InvMM to be a correct measure of memorization, it should be capable of distinguishing highly memorized images from normal ones. To construct a reliable test set, we conduct a nearest neighbor test following previous studies [50, 58, 9] to discover highly memorized images from DM’s training set. In this sense, highly memorized images consist of training images that can appear at sampling time. Specifically, we consider DDPM trained on CIFAR-10, LDMs trained on CelebAHQ and FFHQ, and SD trained on LAION. With their official implementations, only DDPM and Stable Diffusion v1.4 present obvious replications, which are thus used for evaluation. Finally, for DDPM, the 99 discovered training images appearing in 1M random samples are used as positive samples and another 1k different training images are included as negative samples; for SD v1.4, we deem images in the confirmed set as positive and those in the normal set as negative.
@@
Table 1 presents the detection performance assessed via AUC and TPR@1%FPR. For the unconditional DDPM on CIFAR-10, training loss metrics (the first 2 rows) show decent results and the original $\bm{\epsilon}_{0}$-loss works much better than the $\bm{x}_{0}$-loss. The differential metric $M^{\mathrm{LOO}}$ is less effective than the training loss and InvMM. The results indicate that training loss based metrics (including $M^{\mathrm{LOO}}$) are not reliable to signify memorization, as they are not fundamentally commensurable between samples. InvMM measures memorization of each sample by contrasting their sensitive noise distribution to a common standard Gaussian, thus having the same scale.
For SD v1.4, we compare to training loss metrics and the metric proposed by Wen et al. [61]. As the compared methods all require access to image captions, we implement the counterpart for InvMM as well. With access to captions, all the metrics achieve perfect detection performance. However, this does not mean that the compared methods can also be utilized as memorization measures. On the one hand, training captions might not be available in practice, e.g., for non-training images. On the other hand, exploiting training captions for detection is actually to predict whether the captions will replicate the training images, rather than measure the memorization. In contrast, that our measure calculated with prompt optimized simultaneously shows a degree of performance drop, implies that images considered as less memorized according to the behavior of their training captions might be actually memorized more, exposed by potential prompts. In this sense, our measure provides a worst-case estimate of memorization for conditional DMs.
5.4 Influence factors
Both theoretical analysis and empirical investigation point out that dataset complexity [15, 51], dataset size [15, 51], cardinality of latent vector set [53] and data duplication [59] are important factors for memorization. We therefore verify the correctness of our measure by utilizing it to describe tendency of memorization in the change of these factors.
Dataset complexity, size and training epoch. First of all, we simultaneously validate dataset complexity, size and cardinality of latent vector set on CelebAHQ and FFHQ. Following Feng et al. [15], we measure dataset complexity with Intrinsic Dimensionality (ID) [31]. The cardinality of latent vector set can be controlled by training epochs (if in each iteration a fresh random latent vector is picked). However, we argue that previous works [15, 51] obfuscate the effect of dataset size and training epochs, as models are trained for different epochs (probably more) on smaller subsets to the full dataset. Consequently, on smaller subsets, each training image would see different number of latent vectors. Therefore, in our experiment, we decouple these two factors by training for the same epochs as on the full dataset when studying the influence of dataset size.
Figure 5: Dataset complexity & size vs. memorization. All the models are trained for 943 epochs on CelebAHQ and 446 epochs on FFHQ.
Figure 6: Training epoch vs. memorization. Models are trained for multiples of the default epoch (CelebAHQ: 943, FFHQ: 446).
As can be observed in Fig. 5, dataset complexity and size seem not present high correlation with memorization, as indicated by our measure. However, when trained for longer time, the memorization scores gradually decrease, which indicates a large extent of memorization. Therefore, dataset complexity and size are not independent determining factors, and in conclusion, our measure accurately reflects the impact of these three factors on memorization.
Figure 7: Duplication vs. memorization on CelebAHQ-2.5k and FFHQ-6k. CelebAHQ-250 and FFHQ-600 subsets are duplicated 5, 10, 15 and 20 times respectively. Outliers are hidden for clarity.
Figure 8: Deduplication vs. memorization.
Data duplication. We then validate the effect of data duplication in DMs’ training set through our measure. We train a series of LDMs on CelebAHQ-2.5k and FFHQ-6k, with their subsets CelebAHQ-250 and FFHQ-600 duplicated 5, 10, 15 and 20 times. Figure 7 depicts the memorization scores of duplicated images. As can be observed, data duplication largely worsens memorization in DMs and our measure is sensitive to these changes. With increasing duplication times, the larger FFHQ-6k dataset presents a stronger resistance to memorization.
For CIFAR-10 and LAION which originally contains duplicates, we observe the effect of deduplication. We deduplicate CIFAR-10 with SSCD embeddings and DBSCAN clustering [14]. This results in 666 identified clusters, only one image of which is retained for training. The memorization is evaluated on the retained 666 images before and after deduplication. On LAION, as compared to SD v1.4, SD v2.1 deduplicates the training set. The memorization is evaluated on the confirmed and suspicous sets. Figure 8 shows the results, with infinite scores highlighted. As can be observed, InvMM clearly exhibits the benefits that deduplication brings.
@@ 5.5 Comparison to membership
A close topic to memorization is membership inference [48, 8, 13], which aims to predict whether a sample is used for training. The common accepted assumption is that member samples are fit “better” than hold-out samples [13], thus statistics like (variant of) training loss or differential training loss [8] are generally used for membership inference. A comparison to a recent membership inference metric SecMI [13] on FFHQ-6k, as shown in Fig. 9, does present a positive correlation between membership and memorization. As factors that exacerbate memorization changes, both SecMI and InvMM shows better performance.
However, there is still a performance gap between membership score and memorization score, which indicates that certain samples from the hold-out set are memorized more than some member samples. This mainly results from the gap between model’s metric (training loss) and human’s metric (perceptual similarity). Membership inference leverages the model’s metric for prediction, while memorization should resort to perceptual metric, e.g., simulated by SSCD. What level of loss leads to image replication depends both on the model and image itself. Low training loss does not necessarily lead to data replication, but would probably distinguish trained data from never accessed. On the contrary, high training loss indicating non-membership may still be enough for replication. In summary, membership does not deterministically lead to high memorization and vice versa. Our measure differs from membership inference metric by focusing on the perceptual level and adaptively handles the specific loss landscape of different samples.
@@ 6 Conclusion
In this work, we dive into measuring memorization for diffusion models. Comprehensive experiments demonstrate the ability of the proposed measure to correctly describe the extent to which any image is memorized. In short, an image to a higher extent of memorization is able to be inverted to a more standard and diverse latent distribution. However, our method is also limited by its simple modeling of the sensitive latent noise as multivariate normal. During our research, we observed many clues confirming that the true distribution is complex. Future work might consider more advanced methods, e.g., a generative network, to fit the sensitive noise. Our measure may also support more practical applications, such as online sensitive noise filtering and adversarial training to mitigate harmful memorization.
@@ 7.2 Reparameterization
where $\{\bm{g}_{i,j}\},i=1...N,j=1...|\mathcal{V}|$ are i.i.d samples drawn from $\mathrm{Gumbel}(0,1)$, $\tau$ is a temperature factor. When $\tau$ approaches 0, the smoothed sample $\tilde{\bm{\omega}}_{i}$ becomes one-hot.
@@ 8.1 Experiment setting
If not stated otherwise, the hyperparameters follow the default setting listed in Tab. 2. For CelebAHQ and FFHQ, all memorization scores in the plots are evaluated on CelebAHQ-250 and FFHQ-600. All the experiments in this paper are conducted on one NVIDIA A800 GPU.

### v3

@@ 4.4 Text-to-image DM
For text-to-image DM, we model the prompt part as a distribution as well and optimize it simultaneously with the noise distribution. Suppose an input prompt is tokenized into maximum $M$ tokens $\bm{\omega}=[\bm{\omega}_{1},\bm{\omega}_{2},...,\bm{\omega}_{M}]$. Motivated by Guo et al. [20], we learn a $|\mathcal{V}|$-dimensional categorical distribution over the vocabulary $\mathcal{V}$ for each token $\bm{\omega}_{i}$. Let token $\bm{\omega}_{i}$ follows a categorical distribution with probabilities $\bm{\pi}_{i,j}$, where $\sum_{j=1}^{|\mathcal{V}|}\bm{\pi}_{i,j}=1$. $\bm{\pi}_{i,j}$ can be parameterized and optimized.
Because sampling from categorical distribution is non-differentiable, we adopt the Gumbel-Softmax reparameterization [23] to obtain smoothed sample $\tilde{\bm{\omega}}_{i}$. Given the smoothed token $\tilde{\bm{\omega}}_{i}$, it can be converted to token embedding as a linear combination of vocabulary embeddings $e(\tilde{\bm{\omega}}_{i})=\sum_{j=1}^{|\mathcal{V}|}\tilde{\bm{\omega}}_{i,j}e(j)$, where $e(j)$ is the pretrained token embedding of the $j$-th token. A proper choice of the prompt distribution is critical for inversion on text-to-image DMs. We provide a detailed analysis in Appendix Sec. 9.
@@ 5 Experiment
In this section, we conduct a series of experiments from different perspectives to demonstrate the validity of InvMM. First, we conduct controlled studies to verify the effectiveness of the adaptive algorithm (Sec. 5.2). Based on this, we show that InvMM can reasonably elucidate the influence of various factors on memorization (Sec. 5.3) and identify highly memorized images (Sec. 5.4). Finally, we discuss how memorization differs from membership (Sec. 5.5).
@@ 5.1 Experiment setup
Dataset. We use four datasets for evaluation, i.e., CIFAR-10 [31], CelebAHQ [26], FFHQ [27] and LAION-Aesthetics V2 [47]. For subsequent experiments, we construct four randomly sampled subsets of CelebAHQ and FFHQ: CelebAHQ-250, CelebAHQ-2.5k, FFHQ-600 and FFHQ-6k. We ensure that the larger subset includes the smaller subset. The LAION dataset is extremely large. To keep computational costs manageable, we collect a subset of images. We first include 76 heavily memorized images discovered by Webster [58]. We call this set of images the confirmed set. In addition, we collected another 50 images with high duplication rates according to the statistics from Webster et al [60]. This set of suspects, named as the suspicious set, are assured to be disjoint with the confirmed set. Images in the confirmed set can be triggered by their training captions while those in the suspicous set cannot. Finally, another 50 randomly sampled images are used for comparison, referred to as the normal set.
Diffusion Models. Following the implementations of DDPM [22] and Latent Diffusion Model (LDM) [43], we train DDPM on CIFAR-10 and LDM on the subsets of CelebAHQ and FFHQ. Pre-trained LDMs on the whole CelebAHQ and FFHQ are also used. For text-to-image generation, we utilize Stable Diffusion (SD) v1.4, v2.1 and v3.5 trained on LAION.
Inversion. The experiment setting and additional results are deferred to Appendix Sec. 8.
@@ 5.2 Adaptive algorithm
A comparison of the resulting memorization score under different settings can be observed in Fig. 3. Fixing $\lambda$ to 1 cannot successfully invert most images: 92.4% and 99.7% images of CelebAHQ and FFHQ result in infinite memorization scores. In contrast, both $\lambda=0$ and adaptive reweighting successfully invert all images. However, Fixing $\lambda$ to 0 leads to larger resulting memorization scores, which indicates underestimation of the extent of memorization.
Figure 4 showcases the generation results with the learned distributions. With the adaptive algorithm, we successfully perform inversion across various datasets and DMs that the randomly generated samples closely resemble the training images. The results form the premise of InvMM that memorization of samples is measured in the state they are made replicable.
5.3 Influence factors
Both theoretical analysis and empirical investigation point out that dataset complexity [15, 52], dataset size [15, 52], cardinality of latent vector set [54] and data duplication [60] are important factors for memorization. We thus verify the correctness of our measure by utilizing it to describe how memorization trends vary with changes in these factors.
Dataset complexity, size and training epoch. We begin by simultaneously validating dataset complexity, size and cardinality of latent vector set on CelebAHQ and FFHQ. Following Feng et al. [15], we quantify dataset complexity with Intrinsic Dimensionality (ID) [32]. The cardinality of latent vector set can be controlled by adjusting training epochs (if in each iteration a fresh random latent vector is picked). However, we argue that previous works [15, 51] obfuscate the effect of dataset size and training epochs, as models are trained for different epochs (probably more) on smaller subsets to the full dataset. Consequently, on smaller subsets, each training image would see different number of latent vectors. Therefore, in our experiment, we decouple these two factors by training for the same epochs as on the full dataset when studying the influence of dataset size.
Figure 5: Dataset complexity & size vs. memorization. Models are trained for 943 epochs on CelebAHQ and 446 epochs on FFHQ.
Figure 6: Training epoch vs. memorization. Models are trained for multiples of the default epochs.
As can be observed in Fig. 5, dataset complexity and size seem not present high correlation with memorization, as indicated by our measure. When trained for longer time (Fig. 6), each model presents a larger extent of memorization. The results suggest that dataset complexity and size are not independent determining factors, and InvMM accurately reflects the impact of these three factors.
Figure 7: Duplication vs. memorization.
Figure 8: Deduplication vs. memorization.
Data duplication. We then validate the effect of data duplication in DMs’ training set through our measure. We train a series of LDMs on CelebAHQ-2.5k and FFHQ-6k, with their subsets CelebAHQ-250 and FFHQ-600 duplicated 5, 10, 15 and 20 times. Figure 7 depicts the memorization scores of duplicated images. As can be observed, data duplication rapidly worsens memorization in DMs and our measure is sensitive to these changes. With increasing duplication times, the larger FFHQ-6k dataset presents a stronger resistance to memorization.
For CIFAR-10 and LAION which originally contain duplicates, we observe the effect of deduplication. We deduplicate CIFAR-10 with SSCD embeddings and DBSCAN clustering [14]. This results in 666 identified clusters, only one image of which is retained for training. The memorization is evaluated on the retained 666 images before and after deduplication. On LAION, as compared to SD v1.4, SD v2.1 and v3.5 deduplicate the training set. Memorization is evaluated on the confirmed and suspicous sets. Figure 8 shows the results, with infinite scores on the top. As can be observed, InvMM clearly exhibits the benefits that deduplication brings.
@@ 5.4 Detection
$\bm{\epsilon}_{0}$-loss 0.981 0.414 0.955 0.000
$\bm{x}_{0}$-loss 0.180 0.000 0.975 0.342
@@
Wen [61] - - 0.999 0.868
Ren [42] - - 0.975 0.355
InvMM 0.997 0.899 0.974 0.908
Table 1: Detection performance on DDPM and SD v1.4.
Figure 9: (a) Memorization score distribution on three datasets. (b) Adversarial prompt for an example from the suspicious set.
@@
For InvMM to be a correct measure of memorization, it should be capable of distinguishing highly memorized images from normal ones. To construct a reliable test set, we conduct a nearest neighbor test following previous studies [51, 59, 9] to discover highly memorized images from DM’s training set, i.e., highly memorized images consist of training images that can appear at sampling time. Specifically, we consider DDPM trained on CIFAR-10, LDMs trained on CelebAHQ and FFHQ, and SDs trained on LAION. With their official implementations, only DDPM and SD v1.4 present obvious replications, which are thus used for evaluation. For DDPM, the 99 discovered training images appearing in 1M random samples are used as positive samples and another 1k different training images are included as negative samples; for SD v1.4, in accordance with the setting of previous works, we deem images in the confirmed set as positive and those in the suspicious and normal set as negative, considering that captions in the confirmed set can cause replication of their paired images while those in the suspicious and normal set cannot.
@@
Table 1 presents the detection performance assessed via AUC and TPR@1%FPR. For the unconditional DDPM on CIFAR-10, training loss metrics (the first 2 rows) show decent results and the original $\bm{\epsilon}_{0}$-loss works much better than the $\bm{x}_{0}$-loss. The differential metric $M^{\mathrm{LOO}}$ is less effective than the $\bm{\epsilon}_{0}$-loss and InvMM. The results indicate that the compared metrics are not able to quantify memorization, as they are not fundamentally commensurable between samples. InvMM measures memorization of each sample by contrasting its sensitive noise distribution to a common standard Gaussian, thus having the same scale.
For SD v1.4, we compare InvMM against training loss metrics and two state-of-the-art detection metrics [61, 42]. Since all compared approaches require access to image captions, we implement InvMM with fixed prompts during inversion. As shown in Tab. 1, our measure achieves performance comparable to state-of-the-art metrics. The results of compared metrics seem to demonstrate a strong ability to distinguish memorization thus far. However, we argue that the prompt-based task formulation employed in existing studies offers an overly optimistic understanding of memorization. It is because these techniques are actually to predict whether the captions will replicate their paired images, while overlook potential prompts that might cause replication. As observed in Fig. 9(a), when simultaneously inverting the input prompt, our measure reveals that some images in the suspicious set are also memorized to a similar extent as those in the confirmed set (due to duplication). Applying GCG [64] attack on images with the lowest InvMM scores in the suspicious set provides additional evidence, as illustrated in Fig. 9(b): images previously deemed “not memorized” according to their training captions can indeed be replicated by potential adversarial prompts (see Appendix Fig. 13 for more examples). Hence, going beyond the confines of existing research, our measure describes worst-case memorization from an adversarial standpoint, making it more reliable in security-critical scenarios.
@@ 5.5 Comparison to membership
A close topic to memorization is membership inference [49, 8, 13], which reveals the leakage risk of membership information that whether an image is used for training. The common implicit assumption is that member samples are “memorized” more than hold-out samples [13], rendering them distinguishable through statistics like (variants of) training loss or differential loss [8]. Interestingly, the principle of InvMM can provide an explicit clarification of the difference between membership and memorization.
A comparison to a recent membership inference metric SecMI [13] on FFHQ-6k, as shown in Fig. 10, presents a positive correlation between them. As factors that exacerbate memorization changes, both SecMI and InvMM show increasing performance. There is also a performance gap between SecMI and InvMM, indicating that certain samples from the hold-out set are memorized more than some member samples. This mainly results from the gap between model’s metric (training loss) and human’s metric (perceptual similarity). Membership inference leverages the model’s metric for prediction, while memorization should resort to perceptual metric, e.g., simulated by SSCD. What level of loss leads to image replication depends both on the model and image itself. Low training loss does not necessarily lead to data replication, but would probably distinguish training data from never accessed. On the contrary, high training loss signifying non-membership may still be enough for replication. Therefore, membership does not deterministically lead to high memorization and vice versa. InvMM differs from membership metrics by focusing on the perceptual level and adaptively handles the specific loss landscapes of different samples.
@@ 6 Conclusion
In this work, we dive into measuring memorization for diffusion models. Comprehensive experiments demonstrate the ability of the proposed measure to correctly quantify memorization. In short, an image to a higher extent of memorization is able to be inverted to a more standard and diverse latent distribution. However, our method is also limited by its simple modeling of the sensitive latent noise as a multivariate normal. During our research, we observed many clues confirming that the true distribution is complex. Future work might explore more advanced methods, e.g., a generative network, to fit the sensitive noise.
Acknowledgement
This work was partly supported by the National Key Research and Development Program of China under No. 2022YFB3102100, NSFC under No. U244120033, U24A20336, 62172243, 62402425 and 62402418, the China Postdoctoral Science Foundation under No. 2024M762829, the Zhejiang Provincial Natural Science Foundation under No. LD24F020002, the “Pioneer and Leading Goose” R&D Program of Zhejiang under No. 2025C01082, 2025C02033, 2025C02263 and 2025C02034, the Zhejiang Provincial Priority-Funded Postdoctoral Research Project under No. ZJ2024001 and the Key R&D Program of Ningbo under No. 2024Z115.
@@ Acknowledgement
We provide supplementary technical details and experimental results in the following sections:
  • Sec. 7 presents a unified theoretical framework for InvMM within the context of diffusion model, including detailed analyses of noise distribution and prompt distribution reparameterization.
  • Sec. 8 includes additional experimental details and results complementing to the main paper. Specifically, results in Sec. 8.5 provides more evidence to demonstrate the difference between membership and memorization.
  • Sec. 9 gives a comprehensive ablation study on the prompt inversion in text-guided DMs. The experiments clarify the influence of temperature, CFG scale, inversion objective and adaptive algorithm on prompt inversion.
  • Sec. 10 shows generation results using inverted noise and prompts on various datasets and models.
@@ 7.2 Reparameterization
where $\{\bm{g}_{i,j}\},i=1...M,j=1...|\mathcal{V}|$ are i.i.d samples drawn from $\mathrm{Gumbel}(0,1)$, $\tau$ is a temperature factor. When $\tau$ approaches 0, the smoothed sample $\tilde{\bm{\omega}}_{i}$ becomes one-hot.
@@ 8.1 Experiment setting
If not stated otherwise, the hyperparameters follow the default setting listed in Tab. 2, determined by previous investigation [51] and a few case studies. For CelebAHQ and FFHQ, all memorization scores in the plots are evaluated on CelebAHQ-250 and FFHQ-600. SD v3.5 utilizes three text encoders to represent the input prompt: CLIP-L, CLIP-G and T5. It is computationally expensive, so we only invert CLIP-L, with CLIP-G and T5 frozen. All the experiments in this paper are conducted on one NVIDIA A800 GPU.
@@
8.2 Influence factors
Figure 11: Training epoch vs. memorization on CelebAHQ.
Figure 11 shows the influence of training epochs on CelebAHQ. Larger training epochs lead to heavier memorization.